water in a canal 3m wide and 1.2m deep is flowing with a speed of 10km/hr. how much area will it irrigate in 30 minutes, if 8 cm of standing water is required for irrigation

Grade:10

1 Answers

Arun
6 years ago
Depth = 1.2m
width =  3m
flow velocity = 10km/h = 10,000 m/h
time = 30minute = 0.5 hour
standing water required = 8cm = 0.08m
Let the area = A m²
Volume of water will be same
1.2m × 3m × 10,000 m/h × 0.5 h = A × 0.08
⇒ A = 18000/0.08
⇒ A = 225000 m²
Regards
Arun (askIITians forum expert)